# RAFT: Adapting Language Model to Domain Specific RAG

### v1

@@ Abstract
Pretraining Large Language Models (LLMs) on large corpora of textual data is now a standard paradigm. When using these LLMs for many downstream applications, it is common to additionally bake in new knowledge (e.g., time-critical news, or private domain knowledge) into the pretrained model either through RAG-based-prompting, or finetuning. However, the optimal methodology for the model to gain such new knowledge remains an open question. In this paper, we present Retrieval Augmented Fine Tuning (RAFT), a training recipe that improves the model’s ability to answer questions in an "open-book" in-domain setting. In RAFT, given a question, and a set of retrieved documents, we train the model to ignore those documents that don’t help in answering the question, which we call, distractor documents. RAFT accomplishes this by citing verbatim the right sequence from the relevant document that would help answer the question. This coupled with RAFT’s chain-of-thought-style response helps improve the model’s ability to reason. In domain specific RAG, RAFT consistently improves the model’s performance across PubMed, HotpotQA, and Gorilla datasets, presenting a post-training recipe to improve pre-trained LLMs to in-domain RAG. RAFT’s code and demo are open-sourced at https://github.com/ShishirPatil/gorilla
Figure 1: How best to prepare for an Exam?(a) Fine-tuning based approaches implement "studying" by either directly "memorizing" the input documents or answering practice QA without referencing the documents. (b) Alternatively, in-context retrieval methods fail to leverage the learning opportunity afforded by the fixed domain and are equivalent to taking an open-book exam without studying. While these approaches leverage in-domain learning, they fail to prepare for open-book tests. In contrast, our approach (c) RAFT leverages fine-tuning with question-answer pairs while referencing the documents in a simulated imperfect retrieval setting — thereby effectively preparing for the open-book exam setting.
Machine Learning, ICML
[…]
envname-P envname#1
tianjunz@berkeley.edu, shishirpatil@berkeley.edu
UC Berkeley
@@ 1 Introduction
Trained on vast quantities of public data, Large Language Models LLMs have achieved significant advances in a wide range of general knowledge reasoning tasks  (Brown et al., 2020; Wei et al., 2022).
@@
However, increasingly LLMs are being employed in specialized domains to support tasks ranging from code completion for specific software frameworks to question answering on specific document collections (e.g., legal or medical documents). In these settings, general knowledge reasoning is less critical but instead, the primary goal is to maximize accuracy based on a given set of documents. Indeed, adapting LLMs to the specialized domains (e.g., recent news, enterprise private documents, or program resources constructed after the training cutoff) is essential to many emerging applications (Vu et al., 2023; Lazaridou et al., 2022) and is the focus of this work.
@@
This paper studies the following question – How to adapt pre-trained LLMs for Retrieval Augmented Generation (RAG) in specialized domains?
@@
When it comes to adapting LLMs to specialized domains, we consider the following two candidates: in-context learning through Retrieval-Augmented Generation (RAG) and supervised fine-tuning. RAG-based methods allow the LLM to reference the documents when answering questions. However, these methods fail to leverage the learning opportunity afforded by the fixed domain setting and early access to the test documents. Alternatively, supervised fine-tuning offers the opportunity to learn more general patterns in the documents and better align to end tasks and user preferences (Zhou et al., 2023a). However, existing fine-tuning based approaches either fail to leverage the documents at test time (don’t incorporate RAG) or fail to account for the imperfections in the retrieval process during training.
@@
We can draw an analogy to an open-book exam. Existing in-context retrieval methods are equivalent to taking an open-book exam without studying. Alternatively, existing fine-tuning based approaches implement “studying" by either directly “memorizing" (Xiong et al., 2023) the input documents or answering practice questions (Wang et al., 2022) without referencing the documents. While these approaches leverage in-domain learning they fail to prepare for the open-book nature of test setting.
@@
In this paper, we study how to combine supervised fine-tuning (SFT) with retrieval augmented generation (RAG). We propose a novel adaptation strategy – Retrieval-Augmented Fine Tuning (RAFT). RAFT specifically addresses the challenge of fine-tuning LLMs to incorporate domain knowledge while also improving in-domain RAG performance. RAFT aims to not only enable models to learn domain specific knowledge through fine-tuning, but also to ensure robustness against inaccurate retrievals. This is achieved by training the models to understand the dynamics between the question posed (prompt), the domain specific documents retrieved, and the appropriate answer. Going back to our analogy, our approach is analogous to studying for an open-book exam by recognizing relevant, and irrelevant retrieved documents.
@@
In RAFT, we train the model to answer the question (Q) from Document(s) (D*) to generate an answer (A*), where A* includes chain-of-thought (Wei et al., 2022; Anthropic, 2023), and in the presence of distractor documents ($D_{k}$). We explain the methodology in detail in Section 3 and analyze the sensitivity to the number of distractor documents ($k$) at train- and test- time in Section 5. RAFT consistently outperforms Supervised-finetuning both with- and without- RAG across PubMed (Dernoncourt & Lee, 2017), HotpotQA (Yang et al., 2018), and HuggingFace Hub, Torch Hub, and Tensorflow Hub Gorilla datasets (Patil et al., 2023), presenting a novel, yet simple technique to improve pre-trained LLMs for in-domain RAG.
@@ 2 LLMs for Open-Book Exam
To understand our goal better, we expand on our analogy between training an LLM in the real-world setting of preparing for an exam.
@@
A closed book exam often refers to a scenario where the LLMs do not have access to any additional documents or references to answer the questions during the exam. For LLMs, this is equivalent to the scenario, for example, in which the LLM is used as a chatbot. In this scenario, the LLM draws from the knowledge baked in during pre-training and supervised finetuning to respond to the prompt.
@@
In contrast, we liken the open-book exam setting to the scenario in which the LLM can refer to external sources of information (e.g., a website or a book chapter). In such scenarios, typically, the LLM is paired with a retriever which retrieves ‘k’ documents (or specific segments of the document) which are appended to the prompt. It is only through these documents retrieved that the LLM gains access to “new knowledge”. As a result, we argue that the LLM’s performance in these settings, where it is trained as a general-purpose LLM is largely dependent on the quality of the retriever and how accurately the retriever can identify the most relevant piece of information.
Domain Specific Open-Book Exam
In this paper, we focused on a narrower but increasingly popular domain than the general open book exam, called the domain specific open book exam. In domain specific open book exams, we know apriori the domain in which the LLM will be tested – used for inference. The LLM can respond to the prompt using use any and all information from this specific domain, which it has been fine-tuned on. Examples of domain specific examples include enterprise documents, latest news, code repositories belonging to an organization, etc. In all these scenarios, the LLM will be used to respond to the questions, whose answers can be found within a collection of documents (a small practical domain). The retrieval technique itself has little to no impact on the mechanism (though it may impact the accuracy). This paper mainly studies this, domain specific open-book setting and how to adapt a pretrained LLM to this specific domain, including how to make it more robust to a varying number of retrieved documents and distractors.
@@ 3 RAFT
Figure 2: Overview of our RAFT method. The top-left figure depicts our approach of adapting LLMs to reading solution from a set of positive and negative documents in contrast to standard RAG setup where models are trained based on the retriever outputs, which is a mixture of both memorization and reading. At test time, all methods follow the standard RAG setting, provided with a top-k retrieved documents in the context.
In this section, we present RAFT, a novel way of training LLMs for domain specific open-book exams. We first introduce the classical technique of supervised fine-tuning, followed by the key takeaways from our experiments. Then, we introduce RAFT , a modified version of general instruction tuning. Lastly, we provide an overview of the experiments to expect in the later sections.
@@
Consider the supervised fine-tuning (SFT) setting for a Question-Answer dataset. The formulation consists of the Dataset ($D$) from which a set of Question ($Q$) and corresponding answer ($A$) pairs are derived or already available. In the classical SFT setting, the model is trained to improve its ability to answer the questions based on its knowledge - obtained either during pre-training, or during the SFT training phase. The model so trained can also be used at test-time with the Retrieval Augmented Generation (RAG) setting, where additional documents can be introduced in the prompt to help the model answer the question. This can be represented as follows:
  • Train: $\mathbf{Q}$} $\to$ $\mathbf{A}$}
  • 0-shot Inference: $\mathbf{Q}$} $\to$ $\mathbf{A}$}
  • RAG Inference: $\mathbf{Q}$} + $\mathbf{D}$} $\to$ $\mathbf{A}$}
RAFT
Retrieval Aware Fine-Tuning (RAFT), presents a novel recipe to prepare fine-tuning data to tailor the models for domain specific open-book settings, equivalent to in-domain RAG In RAFT, we prepare the training data such that each data point contains a question ($Q$), a set of documents ($D_{k}$), and a corresponding Chain-of-though style answer ($A^{*}$) generated from one of the document ($D^{*}$). We differentiate between two types of documents: ‘oracle’ documents (DD*) i.e. the documents from which the answer to the question can be deduced, and ‘distractor’ documents ($D_{i}$) that do not contain answer-relevant information. As an implementation detail, the ‘oracle’ document doesn’t need to be a single document, but can be more than one document, as is the case in HotpotQA (Yang et al., 2018). Then, for $P$ fraction of the questions ($q_{i}$) in the dataset, we retain the oracle document ($d_{i}^{*}$) along with distractor documents ($d_{k-1}$). For (1P)(1-P) fraction of the questions ($q_{i}$) in the dataset, we include no oracle document and only include distractor documents ($d_{k}$). We then fine-tune the language model using the standard supervised training (SFT) technique, training it to generate answers from the provided documents and questions. Fig. 2 illustrates the high-level design principal for RAFT .
We demonstrate that our approach trains the model to perform better RAG on the set of documents it is trained on i.e., in-domain. By removing the oracle documents in some instances, we are compelling the model to memorize answers instead of deriving them from the context. The training data for RAFT is as follows, and an example of training data can be seen in Fig. 3:
  • $\mathbf{P}$} % of data: $\mathbf{Q}$} + $\mathbf{D}$$^{*}$ + ${\mathbf{D}}_{2}$ + $\dots$ + ${\mathbf{D}}_{k}$ $\to$ $\mathbf{A}$}*
@@
A key factor in enhancing training quality is the generation of a reasoning process, such as Chain-of-Thought, to explain the provided answers.RAFT  approach is similar: we demonstrate that creating a full reasoning chain and in addition, clearly citing sources enhances the model’s accuracy in answering questions. In Fig. 3, we illustrate this set-up. Generating the training data in this fashion, involves presenting the model with a question, context, and verified answers, and then requesting it to form a reasoning chain that appropriately references the original context.
For all the datasets in our experiments, we generate the answers using the technique described above. Note that the Gorilla APIBench dataset, already includes reasoning in the answers. We provide an example of the generation step in Fig. 3, the detailed reasoning answer includes a citation from the original context inside ##begin_quote## and ##end_quote## as well as the detailed explanation on how to reach the conclusion based on the citations. We demonstrate that adding detailed reasoning paragraphs helps boost the model’s performance in our experiment section.
@@
Table 1: RAFT improves RAG performance forall specialized domains: Across PubMed, HotpotQA, HuggingFace, Torch Hub, and Tensorflow Hub, we see that domain specific Finetuning improves significantly of the performance of the base model, but RAFT  consistently outperforms the existing domain specific finetuning method with or without RAG. This suggests the need to train the model with context. We compare our model with LLaMA finetuning receipes, and provide GPT-3.5 for reference.
PubMed HotpotQA HuggingFace Torch Hub TensorFlow Hub
@@ 4 Evaluation
We design our experiments to study how well RAFT  performs compared to various baselines. We find that the RAFT-7B model (a finetuned version of LlaMA-2) is better at reading and extracting information from in-domain documents, than domain specific finetuned model, and general-purpose model with RAG. As an ablation, we also demonstrate how important it is for the model to learn with Chain-of-Thought responses. In this section, we will first introduce all the datasets we used in the experiments, then all the baseline model/fine-tuning techniques that we benchmark against.
4.1 Datasets
In our experiments, we use the following datasets to evaluate our model and all baselines. We selected these datasets to represent both popular and diverse domains including Wikipedia, Coding/API documents, and question-answering on medical documents.
  • Natural Questions (NQ) (Kwiatkowski et al., 2019), Trivia QA (Joshi et al., 2017) and HotpotQA (Yang et al., 2018) are the open-domain question-answers based on Wikipedia, mainly focused on common knowledge (e.g., movies, sports, etc).
  • HuggingFace, Torch Hub, and TensorFlow Hub are from the APIBench (Patil et al., 2023) proposed in the Gorilla paper. These benchmarks measure how to generate the correct, functional, and executable API calls based on the documentation.
  • PubMed QA (Jin et al., 2019) is a question-answering dataset tailored only for biomedical-research question-answering. It mainly focuses on answering medical and biology questions based on a given set of documents.
Note that the first category of dataset (NQ, Trivia QA, and HotpotQA) is a relatively general domain whereas the latter two domains are on very domain specific documents.
@@ 4.1 Datasets
  • LlaMA2-7B-chat model with RAG (Llama2 + RAG): similar to the previous setting, except here we include reference documents. This is a popular technique when dealing with domain specific QA tasks.
  • domain specific Finetuning with 0-shot prompting (DSF): Performing standard supervised finetuning, without documents in context. We find that it mostly useful to align the answering style of the model as well as get familiar with the domain context.
  • domain specific Finetuning with RAG (DSF + RAG): Equip a domain specific finetuned model with external knowledge using RAG. So, for the “knowledge” the model does not know, it can still refer to the context.

### v2

# envname-P envname#1

@@ Abstract
Pretraining Large Language Models (LLMs) on large corpora of textual data is now a standard paradigm. When using these LLMs for many downstream applications, it is common to additionally incorporate new information into the pretrained model either through RAG-based-prompting, or finetuning. However, the best methodology to incorporate information remains an open question. In this paper, we present Retrieval Augmented Fine Tuning (RAFT), a training recipe which improves the model’s ability to answer questions in "open-book" in-domain settings. In training RAFT, given a question, and a set of retrieved documents, we train the model to ignore those documents that don’t help in answering the question, which we call, distractor documents. RAFT accomplishes this by citing verbatim the right sequence from the relevant document to help answer the question. This coupled with RAFT’s chain-of-thought-style response helps improve the model’s ability to reason. In domain specific RAG, RAFT consistently improves the model’s performance across PubMed, HotpotQA, and Gorilla datasets, presenting a post-training recipe to improve pre-trained LLMs to in-domain RAG.
@@ 1 Introduction
Trained on vast quantities of public data, Large Language Models LLMs have achieved significant advances in a wide range of general knowledge reasoning tasks  Brown et al. (2020); Wei et al. (2022). However, increasingly LLMs are being employed in specialized domains to support tasks ranging from code completion for specific software frameworks to question answering on specific document collections (e.g., legal or medical documents). In these settings, general knowledge reasoning is less critical and instead the primary goal is to maximize accuracy based on a given set of documents. Indeed, adapting LLMs to the specialized domains (e.g., recent news, enterprise private documents, or program resources constructed after the training cutoff) is essential to many emerging applications (Vu et al., 2023; Lazaridou et al., 2022) and is the focus of this work.
@@
This paper studies the following question – How do we adapt pre-trained LLMs for Retrieval Augmented Generation (RAG) in specialized domains?
@@
When it comes to adapting LLMs to specialized domains, we consider the following two candidates: in-context learning through Retrieval-Augmented Generation (RAG) and supervised fine-tuning. RAG based methods allow the LLM to reference the documents when answering questions. However, RAG based in-context learning methods fail to leverage the learning opportunity afforded by the fixed domain setting and early access to the test documents. Alternatively, supervised fine-tuning offers the opportunity to learn more general patterns in the documents and better align to end tasks and user preferences Zhou et al. (2023). However, existing fine-tuning based approaches either fail to leverage the documents at test time (don’t incorporate RAG) or fail to account for the imperfections in retrieval process during training.
@@
We can draw an analogy to an open-book exam. Existing in-context retrieval methods are equivalent to taking an open-book exam without studying. Alternatively, existing fine-tuning based approaches implement “studying" by either directly “memorizing" Xiong et al. (2023) the input documents or answering practice questions Wang et al. (2022) without referencing the documents. While these approaches leverage in-domain learning they fail to prepare for the open-book nature of the test setting.
Figure 1: How best to prepare for an Exam?(a) Fine-tuning based approaches implement "studying" by either directly "memorizing" the input documents or answering practice QA without referencing the documents. (b) Alternatively, in-context retrieval methods fail to leverage the learning opportunity afforded by the fixed domain and are equivalent to taking an open-book exam without studying. In contrast, our approach (c) RAFT leverages fine-tuning with question-answer pairs while referencing the documents in a simulated imperfect retrieval setting — thereby effectively preparing for the open-book exam setting.
@@
In this paper, we study how to combine instruction fine-tuning (IFT) with retrieval augmented generation (RAG). We propose a novel adaptation strategy – Retrieval-Augmented Fine Tuning (RAFT). RAFT specifically addresses the challenge of fine-tuning LLMs to both incorporate domain knowledge while also improving in-domain RAG performance. RAFT aims to not only enable models to learn domain-specific knowledge through fine-tuning, but also to ensure robustness against distracting retrieved information. This is achieved by training the models to understand the dynamics between the question (prompt), the domain-specific documents retrieved, and the right answer. Going back to our analogy to the open book exam, our approach is analogous to studying for an open-book exam by recognizing relevant, and irrelevant retrieved documents.
@@
In RAFT, we train the model to answer the question (Q) from Document(s) (D*) to generate answer (A*), where A* includes chain-of-thought reasoning Wei et al. (2022); Anthropic (2023), and in the presence of distractor documents ($D_{k}$). We explain the methodology in Section 3 and analyze the sensitivity to the number of distractor documents ($k$) at train- and test- time in Section 5. RAFT consistently outperforms Supervised-finetuning both with- and without- RAG across PubMed Dernoncourt & Lee (2017), HotPot QA Yang et al. (2018), and HuggingFace Hub, Torch Hub, and Tensorflow Hub Gorilla datasets Patil et al. (2023), presenting a novel, yet simple technique to improve pre-trained LLMs for in-domain RAG. Our code is available at https://github.com/ShishirPatil/gorilla.
@@ 2 LLMs for Open-Book Exam
To understand our goal better, we expand on our analogy between training an LLM with the real-world setting of prepararing for an exam.
@@
A closed book exam often refers to the scenario where the LLMs do not have access to any additional documents or references to answer the questions during the exam. For LLMs, this is equivalent to the scenario, for example, in which the LLM is used as a chatbot. In this scenario the LLM draws from the knowledge baked in during pre-training and supervised-finetuning to respond to the users’ prompt.
@@
In contrast, we liken the open-book exam setting to the scenario in which the LLM can refer to external sources of information (e.g., a website or a book chapter). In such scenarios, typically, the LLM is paired with retriever which retrieves ‘k’ documents (or specific segments of the document) which are appended to the users’ prompt. It is only through these documents retrieved that the LLM gains access to “domain-specific information”. As a result, we argue that the LLM’s performance in these settings, where it is trained as a general-purpose LLM is largely dependent on the quality of the retriever and how accurately the retriever can identify the most relevant piece of information.
Domain-Specific Open-Book Exam
In this paper, we focus on the narrower but increasingly popular domain than the general open book exam, which we call the domain-specific open-book exam. Here, we know apriori the domain in which the LLM will be tested. The LLM can respond to the users’ prompt using use any and all information from this specific domain, which it has been fine-tuned on. Examples of domain specific examples include enterprise documents, code repositories belonging to an organization, etc. In all these scenarios, the LLM will be used to respond to the questions, whose answers can be found within a collection of documents. The retrieval technique itself has little to no-impact on the mechanism (though it may impact the accuracy). This paper studies the domain-specific open-book setting and how to adapt a pretrained LLM to this specific domain, including how to make it more robust to a varying number of retrieved documents and distractors.
@@ 3 RAFT
Figure 2: Overview of our RAFT method. The top-left figure depicts our approach of adapting LLMs to reading solution from a set of positive and distractor documents in contrast to standard RAG setup where models are trained based on the retriever outputs, which is a mixture of both memorization and reading. At test time, all methods follow the standard RAG setting, provided with a top-k retrieved documents in the context.
In this section, we present RAFT, a novel way of training LLMs for domain-specific open-book exams. We first introduce the classical technique of supervised fine-tuning, followed with the key takeaways from our experiments. Then, we introduce RAFT , a modified version of general instruction tuning. Lastly, we provide an overview of the experiments to expect in the later sections.
@@
Consider the supervised fine-tuning (SFT) setting for a Question-Answer dataset. The formulation consists of the Dataset ($D$) from which a set of Question ($Q$) and corresponding answer ($A$) pairs are derived or already available. In the classical SFT setting, the model is trained to improve it’s ability to answer the questions based on it’s knowledge - obtained either during pre-training, or during the SFT training phase. The model so trained can also be used at test-time with Retrieval Augmented Generation (RAG) setting, where additional documents can be introduced in the prompt to help the model answer the question. This can be represented as follows:
{Train: $\mathbf{Q}$} $\to$ $\mathbf{A}$}}, {0-shot Inference: $\mathbf{Q}$} $\to$ $\mathbf{A}$}}, {RAG Inference: $\mathbf{Q}$} + $\mathbf{D}$} $\to$ $\mathbf{A}$}}
RAFT: Retrieval Augmented Fine-Tuning (RAFT), presents a novel recipe to prepare fine-tuning data to tailor the models for domain-specific open-book setting, equivalent to in-domain RAG In RAFT, we prepare the training data such that each data point contains a question ($Q$), a set of documents ($D_{k}$), and a corresponding Chain-of-though style answer ($A^{*}$) generated from one of the document ($D^{*}$). We differentiate between two types of documents: ‘golden’ documents (DD*) i.e. the documents from which the answer to the question can be deduced, and ‘distractor’ documents ($D_{i}$) that do not contain answer-relevant information. As an implementation detail, the ‘golden’ document doesn’t need to be a single document, but can be more than one document, as is the case in HotpotQA Yang et al. (2018). Then, for $P$ fraction of the questions ($q_{i}$) in the dataset, we retain the golden document ($d_{i}^{*}$) along with distractor documents ($d_{k-1}$). For (1P)(1-P) fraction of the questions ($q_{i}$) in the dataset, we include no golden document and only include distractor documents ($d_{k}$). We then fine-tune the language model using standard supervised training (SFT) technique, training it to generate answers from the provided documents and question. Fig. 2 illustrates the high-level design principal for RAFT .
We demonstrate that our RAG approach trains the model to perform better RAG on the set of documents it is trained on i.e., in-domain. By removing the golden documents in some instances, we are compelling the model to memorize answers instead of deriving them from the context. The training data for RAFT is as follows, and an example training data can be seen in Fig. 3:
$\mathbf{P}$} % of data: $\mathbf{Q}$} + $\mathbf{D}$$^{*}$ + ${\mathbf{D}}_{1}$ + ${\mathbf{D}}_{2}$ + $\dots$ + ${\mathbf{D}}_{k}$ $\to$ $\mathbf{A}$}*
@@
A key factor in enhancing training quality is the generation of a reasoning process, such as Chain-of-Thought, to explain the provided answers. RAFT  approach is similar: we demonstrate that creating a full reasoning chain and in-addition, clearly citing sources enhances the model’s accuracy in answering questions. In Fig. 3, we illustrate this set-up. Generating the training data in this fashion, involves presenting the model with a question, context, and verified answers, and then requesting it to form a reasoning chain that appropriately references the original context.
For all the datasets in our experiments, we generate the answers using the technique described above. Note that the Gorilla APIBench dataset, already includes reasoning in the answers. We provide an example of the generation step in Fig. 3, the detailed reasoning answer includes a citation from the original context inside ##begin_quote## and ##end_quote## as well as the detailed explanation on how to reach the conclusion based on the citations. We demonstrate that adding detailed reasoning paragraphs can help boost the model’s performance in our experiment section.
@@
Table 1: RAFT improves RAG performance for all specialized domains: Across PubMed, HotPot, HuggingFace, Torch Hub, and Tensorflow Hub, we see that Domain-specific Finetuning improves significantly of the performance of the base model, RAFT  consistently outperforms the existing domain-specific finetuning method with or without RAG. This suggests the need to train the model with context. We compare our model with LLaMA finetuning receipes, and provide GPT-3.5 for reference.
PubMed HotPot HuggingFace Torch Hub TensorFlow
@@ 4 Evaluation
We design our experiments to study how well RAFT  performs compared to various baselines. We find that the RAFT-7B model (a finetuned version of LlaMA-2) is better at reading and extracting information from in-domain documents, than domain-specific finetuned model, and general-purpose model with RAG. As an ablation, we also demonstrate how important it is for the model to learn with Chain-of-Thought responses. In this section, we will first introduce all the datasets we used in the experiments, then all the baseline model/fine-tuning techniques that we benchmark against.
Datasets
In our experiments, we use the following datasets to evaluate our model and all baselines. We selected these datasets to represent both popular and diverse domains including Wikipedia, Coding/API documents, and question-answering on medical documents. Natural Questions (NQ) Kwiatkowski et al. (2019), Trivia QA Joshi et al. (2017) and HotpotQA Yang et al. (2018) are the open-domain question-answers based on Wikipedia, mainly focused on common knowledge (e.g., movies, sports, etc). HuggingFace, Torch Hub, and TensorFlow Hub are from the APIBench Patil et al. (2023) proposed in the Gorilla paper. These benchmarks measure how to generate the correct, functional, and executable API calls based on the documentation. PubMed QA Jin et al. (2019) is a question-answering dataset tailored only for biomedical-research question-answering. It mainly focuses on answering medical and biology questions based on a given set of documents. We would like to highlight that (NQ, Trivia QA, and HotpotQA) are relatively general domain whereas the latter two domains are on domain-specific documents.
@@
  • LlaMA2-7B-chat model with RAG (Llama2 + RAG): similar to the previous setting, except here we include reference documents. This is a popular technique when dealing with domain-specific QA tasks.
  • Domain-Specific Finetuning with 0-shot prompting (DSF): Standard supervised-finetuning, without documents in context. We find that its mostly useful to align the answering style of the model as well as get familiar with the domain context.
  • Domain-Specific Finetuning with RAG (DSF + RAG): Equip a domain-specific finetuned-model with external knowledge using RAG. So, for the “knowledge” the model does not know, it can still refer to the context.